Project

General

Profile

Ops/Comm Scheduling meeting

Location : CR21 @ 15:30 pm on September 27th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Hall work

  • DC report
  • Magnet soak test
  • Decay solenoid leak status
  • Water status

Operations

Notes

Magnet soak test

Expert led run on Wednesday pm, identified issues in DAQ, water and run control. Ramped up magnets on Thursday by 11 pm, with overnight cover. Many alarms, but no emergencies. Magnets were calm. Shifter training finished on Friday morning and data taking commenced around lunch time. Smooth data taking until Saturday evening when T01 spontaneously ramped down. Experts were notified and decision was taken to ramp down the magnets. Shifts on Sunday and Monday were spent fixing bugs in the run control and tracker DAQ. ISIS went off on Monday night and is likely to be off until Thursday am at the earliest. See MOM report for summary of data taking and issues.

Decay solenoid

HPC came in Tuesday morning - they repaired seals serviced in the summer; Vicky has restarted and cleaned the system, watching contamination to expose any continuing leak. Signs are hopeful.

Water Status

<Rogers was late due to overrunning MIPO - craig will report>

Detectors

Tracker status (Ed Overton)
  • No report on data quality (Melissa)
  • Started DAQ investigation - VME crate problems are back; tried to take a calibration run this afternoon at 3 pm; looks like communication has failed between cards in the downstream tracker. Will talk with Craig after the meeting. Aim is to poke the system and diagnose which card is causing the problem.
  • Risk for data taking this week; TOF calibration is a fall back position (does not require tracker). Detector alignment is at risk.
KL Status
  • All okay
TOF status
  • Note TOF01/12 plots look interesting - need calibration
  • Note dead/dying channels in TOF0 and TOF2
EMR status
  • Note beam misalignment; 100 mm! SSD misalignment?
  • Known dead channels in MAPMTs
Ckov status
  • All okay

AOB

Alan Young will be doing shifter cover on Monday

tof_quality.pdf (37.4 KB) Wilbur, Scott, 27 September 2016 13:24

8281_emr_status.pdf (297 KB) Drielsma, Fran├žois, 27 September 2016 14:15

KL.pdf (25.2 KB) Bogomilov, Mariyan, 27 September 2016 14:18